### v1

# Learning phases with Quantum Monte Carlo simulation cell

Abstract
We propose a new machine learning input data type, “spin-opstring”, derived from Stochastic Series Expansion Quantum Monte Carlo (QMC) simulations. It offers a compact, memory-efficient representation of QMC simulation cells, combining the initial state with an operator string that encodes the state’s evolution through imaginary time. Using supervised machine learning on two models, we demonstrate the input’s effectiveness in capturing both conventional and topological phase transitions. Additionally, we conduct a regression task to predict superfluid density, which reflects non-local properties of the quantum system, and achieve good accuracy. These results validate the spin-opstring as an effective input for machine learning applications.
I Introduction
In the past decade, due to their ability to analyze complex, high-dimensional data and uncover patterns that are challenging to detect using traditional methods, neural networks and machine learning (ML) techniques have become an integral part of condensed matter physics research (for a review, see for e.g. [1, 2, 3]). In particular, classifying phases and learning information about them, using supervised and unsupervised methods, have been a central focus of the ML approach [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. In this category of ML, various kinds of input data, such as spin configuration, correlation functions, entanglement spectra, winding numbers etc. are used to train the model [10, 16, 7, 11, 9, 17, 18]. In line with the philosophy of artificial intelligence, it is always better to apply as little preprocessing as possible to the training data such that the ML model is able to autonomously identify the relevant features.
There have been proposals to use the raw $(d+1)$-dimensional configuration space generated by quantum Monte Carlo (QMC) methods to train ML models for a $d$-dimensional system, where the additional dimension corresponds to the imaginary time $\beta=1/T$ (with $T$ being the temperature). For increasing system sizes and low temperatures, which is necessary to study quantum phase transitions, the imaginary time direction increases rapidly leading to large simulation cells, which is too large to be effectively handled by machine learning algorithms. In Ref. [10, 19], the auxiliary field configurations obtained from determinant QMC simulations was utilized successfully as ML input, however, for only small systems at high temperatures. To address the challenge of large simulation cells, the authors of Ref. [20] proposed a compression scheme for the $(d+1)$-dimensional space-time configurations generated by Stochastic Series Expansion QMC, to reduce their size to a manageable level and successfully used them as input for a convolutional neural network (CNN).
The Stochastic Series Expansion (SSE) algorithm, developed by Sandvik [21, 22, 23], is a powerful QMC method particularly well-suited for studying quantum spin systems and bosonic models. It is based on a Taylor series expansion of the partition function, the terms of which are then sampled stochastically. Using two update procedures, namely the diagonal and off-diagonal operator update, the SSE method generates simulation cells which can be represented by an allowed arrangement of operators (or operator strings) acting on an initial state. The operator string uniquely defines how the state propagates through the imaginary time and returns back to itself. The periodicity of the state in the imaginary time direction reflects the trace operation in the partition function.
During an SSE QMC simulation, the full set of propagated states is not stored on the computer memory. Instead, only one propagated state and the operator string are stored for each QMC update, from which the remaining states can be generated on the fly. This approach substantially reduces memory requirements compared to other world-line QMC methods [24], where entire space-time configurations are stored. Leveraging this unique feature of SSE, we propose using the initial state along with the operator string, which we call “spin-opstring”, as input data for ML models. In comparison to previous studies that have used QMC simulation cells as input, our input preserves all the information contained in a simulation cell while being memory-efficient.
To demonstrate the effectiveness of the spin-opstring ML input, we applied supervised machine learning techniques to two different models. The first model is the bosonic $t_{2}-V_{1}$ model, which, as reported in Ref. [25], is the simplest model to realize a checkerboard supersolid phase on a two-dimensional square lattice. Using supervised ML with labeled spin-opstring input data generated by SSE QMC, we successfully reproduced the phase transitions of this model. Additionally, we performed a regression task with the ML model to predict the superfluid density, which is in close agreement to the QMC results. To further evaluate the performance of the spin-opstring data in the context of topological phase transitions, we applied supervised ML to the one-dimensional Su-Schrieffer-Heeger (SSH) model [26]. We found that despite the absence of an obvious order parameter, the spin-opstring data is able to capture the transition correctly and predicts the transition point accurately.
The paper is organized in the following manner. In Sec. II we give a brief description of the SSE QMC algorithm and describe the spin-opstring input data. In Sec. III, we outline the input data format for machine learning, followed by a discussion on the neural network architecture and its training procedure. Next, we present the results obtained from the ML model in Sec. IV, and conclude with a summary of our findings in Sec. V.
II Proposal
In quantum statistical mechanics, the primary quantity of interest is the expectation value of any operator $O$ in a system governed by a Hamiltonian $H$ at temperature $T$,
$\langle O\rangle=\frac{1}{Z}\mathrm{Tr}\{Oe^{-\beta H}\},Z=\mathrm{Tr}\{e^{-\beta H}\},$ (1)
where $\beta=1/T$ and all fundamental constants are set to unity. In the presence of non-commuting operators in a Hamiltonian, the calculation of the exponential function $e^{-\beta H}$ in the partition function becomes non-trivial. The starting point of the Stochastic Series Expansion (SSE) method is the expansion of the exponential operator in the partition function in a Taylor series [21],
$Z=\sum_{\alpha}\sum\limits_{n=0}^{\infty}\frac{\beta^{n}}{n!}\langle\alpha|(-H)^{n}|\alpha\rangle.$ (2)
Here, the trace is replaced by a sum over a complete set of basis states $\{|\alpha\rangle\}$ of the Hamiltonian in question. Next, the Hamiltonian is decomposed into classes of elementary interactions as,
$H=-\sum_{a,b}H_{a,b},$ (3)
where the index $a$ denotes the operator class, such as diagonal $(a=1)$ or off-diagonal $(a=2)$ in the chosen basis, while $b$ represents the bond index, which may correspond to nearest-neighbor, next-nearest-neighbor, or any other bond type. With this decomposition, the $n$th power of the Hamiltonian in the partition function can be written as a sum over all possible products, or “strings”, of the operators $H_{a,b}$ (denoted by $S_{n})$ :
$(-H)^{n}=\sum_{S_{n}}\prod\limits_{p=1}^{n}H_{a(p),b(p)}.$ (4)
This results in the partition function
$Z=\sum_{\alpha}\sum\limits_{n=0}^{\infty}\sum_{S_{n}}\frac{\beta^{n}}{n!}\langle\alpha|\prod\limits_{p=1}^{n}H_{a(p),b(p)}|\alpha\rangle.$ (5)
In the above expression, the summation over $n$ varies from 00 to $\infty$, resulting in operator strings or “opstrings” $S_{n}$ of varying length, which are numerically inconvenient to handle. To overcome this difficulty the Taylor expansion is truncated at some finite cutoff $M$, the value of which is dynamically chosen in the algorithm in such a way that the exponentially small truncation errors can be neglected. For the convenience in computation, $(M-n)$ identity operators, $H_{0,0}=I$, are added to a string of $n$ non-identity operators so that the length of all opstrings become equal. Now, $(M-n)$ identity operators can be inserted among the $M$ total operators in $\binom{M}{M-n}=\frac{M!}{(M-n)!n!}$ ways. To avoid over-counting due to the identical nature of the identity operators, we can divide the partition function by $\binom{M}{M-n}$, leading to the expression,
$Z=\sum_{\alpha}\sum_{S_{M}}\frac{\beta^{n}(M-n)!}{M!}\langle\alpha|\prod\limits_{i=1}^{M}H_{a(i),b(i)}|\alpha\rangle.$ (6)
Now, the terms in the above expression can be used as relative probabilities for Monte-Carlo sampling, provided they are positive. While the positive-definiteness of the diagonal terms are always guaranteed, which can be achieved by simply adding a constant, the same condition for the off-diagonal terms requires the lattice to be bipartite.
Figure 1: Schematic diagram of the simulation cell of SSE QMC algorithm for a spin-$1/2$ system with six sites. Here $i$ signifies the lattice site number and $\sigma(i)$ represents the initial/final spin state. At each level $p$ of the simulation cell, the spin state is depicted by open (+1)/filled (-1) circles. The diagonal or off-diagonal operator at each level acting on a bond is represented by open or filled bars respectively. The information in the operator string is compactly codified in $S(p)$ ($a(p)$ and $b(p)$ are defined in the text). The “spin-opstring” for the above simulation cell is constructed by combining $\sigma(i)$ and $S(p)$ together.
The update procedures in SSE QMC algorithm ensure sampling of states with non-zero contributions to the trace while preserving ergodicity. It is achieved by employing two types of updates diagonal update and off-diagonal update. The diagonal update adjusts the number of non-identity Hamiltonian operators $n$ in the operator string $S_{M}$, by replacing diagonal operators with identity ones and vice-versa. On the other hand, off-diagonal update replaces certain diagonal operators with off-diagonal ones and vice-versa, without altering the total number of operators $n$ in the operator strings. For a pedagogical introduction to the various SSE update procedures including recent developments, see [27].
Fig. 1 depicts an example of a typical configuration or simulation cell generated by SSE QMC for a spin chain of $6$ sites. Here, the propagated states at a level $p$ (with $p=0,1,\cdots,M-1$),
$|\alpha(p)\rangle=\prod\limits_{i=1}^{p}H_{a(i),b(i)}|\alpha\rangle,|\alpha(0)\rangle=|\alpha\rangle,$ (7)
are represented by solid and open circles corresponding to $\uparrow$ and $\downarrow$ spins, respectively, at the lattice sites. During an SSE simulation, the operator string $S_{M}$ along with one of the states $|\alpha(p)\rangle$ are stored. Therefore, in contrast to worldline simulations, where the whole space-time spin lattice is stored, the memory requirements for SSE are considerably smaller.
Taking advantage of this specific feature of SSE, in this paper, we propose the initial state $|\alpha(0)\rangle$ along with the operator string $S_{M}$ as the input data for ML models. In contrast to previous studies, where simulation cells has been used as input of ML models, our input retains the entire information stored in a simulation cell, without being memory-extensive.
In SSE simulation, the $\uparrow$ and $\downarrow$ spins are usually stored as $+1$ or $-1$, represented by $\sigma(i)$ in Fig. 1, where $i$ indicates the site number. Moreover, in Fig. 1 the open and solid bars indicate diagonal and off-diagonal operators, respectively, while the absence of a bar between states correspond to a fill-in identity operator. The bond number on which an operator acts at level $p$ is denoted by $b(p)$. If a diagonal operator is present at level $p$, a value $1$ is assigned to the variable $a(p)$, whereas a value $2$ is assigned for an off-diagonal operator. Using this convention, an operator string $S_{M}$ can be constructed such that the diagonal and off-diagonal operators are encoded as even and odd integers, respectively, following the rule $S(p)=2b(p)+a(p)-1$, whereas for identity operators $S(p)$ is assigned to a value 00. We append this operator string $S_{M}$ after the initial spin configuration and prepare the one-dimensional “spin-opstring” input data.
For a system with $N_{s}$ number of sites and an operator string of length M, the storage requirements for the total simulation cell is $N_{s}(1+M)$ memory units, whereas for the spin-opstring data it is just $(N_{s}+M)$. As an example for a $12\times 12$ square lattice ($N_{s}=144$) with a typical $M=25000$, the spin-opstring requires about $140$ times less memory than the whole simulation cell.
III Method
Before feeding the spin-opstring data to the machine, we preprocess the data by applying the following three steps. Firstly, the initial state $|\alpha\rangle$ is mapped from the spin basis to the occupation basis, such that $\uparrow$ spin at site $i$ ($\sigma(i)=1$) maps to the number operator $n_{i}=1$ and $\downarrow$ spin ($\sigma(i)=-1$) maps to $n_{i}=0$. Secondly, the operator string is normalized between values 00 and $1$, by dividing the operator string elements by the maximum possible value, which is $12N_{b}+$ 1, where $N_{b}$ is the total number of bonds in the system. Lastly, since the length of the operator string varies is SSE QMC simulations at different regions of the phase diagram, we determine the maximum length of the training as well as prediction data and apply a padding of zeroes at the end of the operator strings as required to equalize the length of the data. As a zero in the operator string signifies the identity operator, padding with zeroes at the end does not change the physical properties of the system encoded in the data.
To reduce the correlation in the data, we collected spin-opstring data separated by several QMC steps. For example, the training data for phase classification in the $t_{2}-V_{1}$ model at each parameter value consists of 200 spin-opstring data picked from a QMC run of length $10^{5}$ MC steps. The phase classification tasks are achieved with the data being labeled by one-hot encoded vectors, while for regression, we provide the QMC expectation value of the observable as labels. During prediction, for each value of the tuning parameter, the ML model’s output is averaged over multiple spin-opstring data sets obtained from a QMC run at the corresponding parameter value, for both classification and regression tasks. This mimics the idea of an expectation value of an observable for the case of regression. In the classification task we predict the phase boundaries by the point of maximum confusion in the model output, i.e., where the outputs of two softmax neurons intersect.
To analyze the one-dimensional spin-opstring data, we have used the usual feed-forward neural network architecture with fully-connected layers and/or one-dimensional convolutional neural network (CNN) layers. Each of these layers consists of an initial fully connected or convolution step, followed by batch normalization and a rectified linear unit (ReLU) activation function. For the final layer, a fully connected layer is used with a softmax activation function for classification tasks, and a ReLU activation for regression tasks, without batch normalization in both cases. During training, we have minimized the categorical cross-entropy loss function for classification tasks and the mean squared error for regression. The stochastic gradient descent learning process has been optimized by using the Adam algorithm [28]. The code has been implemented in Python by using the Keras library [29]. The ML models used in our study are schematically represented in Table 1 and are further described in detail below.
For the phase classification in the two-dimensional $t_{2}-V_{1}$ model, we have used a ML model with three fully-connected layers: the first input layer consisting of a number of neurons equal to the spin-opstring data length, followed by a hidden layer of 32 neurons, and finally an output layer consisting of a few neurons, depending on the number of phases, with softmax activation. In contrast, for regression, the ML model consists of two fully-connected hidden layers, each with $64$ neurons and the final layer has a single neuron with a ReLU activation function.
Classification model Regression model
FC $n=\mathrm{input}$ BatchNorm ReLU Conv1D $f_{n}=16$, $f_{s}=9$ BatchNorm ReLU FC $n=\mathrm{input}$ BatchNorm ReLU
Maxpool $s=2$
FC $n=32$ BatchNorm ReLU Conv1D $f_{n}=32$, $f_{s}=5$ BatchNorm ReLU FC $n=64$ BatchNorm ReLU
Maxpool $s=2$
Conv1D $f_{n}=64$, $f_{s}=5$ BatchNorm ReLU FC $n=64$ BatchNorm ReLU
Maxpool $s=2$
Conv1D $f_{n}=128$, $f_{s}=5$ BatchNorm ReLU
Maxpool $s=2$
FC $n=2$ or $3$ Softmax FC $n=1$ ReLU
Table 1: Architectural setup of ML models used for classification and regression tasks. The abbreviations used above are as follows: FC: fully connected, $n$ : neuron number, $f_{n}$ : filter number, $f_{s}$ : filter size and $s$ : stride. In the first layers of the fully connected architectures, the number of neurons is equal to the size of the input spin-opstring data.
For phase boundary detection in the 1D SSH model, we have used a CNN architecture. After the input layer, this architecture includes $4$ hidden CNN layers with $16,32,64$ and $128$ filters of sizes $9,5,5$ and $5$, respectively. Each CNN layer is followed by a MaxPool layer with a stride of $2$, which reduces the length of the 1D data by half. The output of the last CNN layer is then passed to a fully-connected layer with two neurons, corresponding to the two phases of the SSH model. For each of the models discussed above, we present our results obtained as an average of the predictions from 100 independently trained instances of the same model.
IV Applications
IV.1 Two-dimensional $t_{2}-V_{1}$ model
Figure 2: Phase diagram of the $t_{2}-V_{1}$ model in Eq. 8 in terms of particle-density $\rho$ and NN repulsion $V_{1}$, where the NNN hopping $t_{2}$ is set to unity. The dashed and solid blue lines represent mixed and continuous phase transitions, respectively, between the SF and cSS phases. The green filled circle denotes the tricritical point separating the mixed transition line from the continuous one. The shaded vertical and horizontal bands indicate the lines where we apply our ML model to predict the phase transitions.
We first apply our proposal to the two-dimensional (2D) $t_{2}-V_{1}$ model which belongs to the family of extended Bose-Hubbard models in the hardcore limit. This simple model displays a rich phase diagram and is, in fact, the minimum model to realize lattice supersolidity of the checkerboard-type. Interestingly, the system also exhibits a line of mixed-order phase transitions with varying particle density which displays the extreme Thouless effect at half-filling. In the 2D $t_{2}-V_{1}$ model, a system of hard-core bosons (HCBs) is studied on a periodic square lattice described by the Hamiltonian,
$H=-t_{2}\sum_{\langle\langle i,j\rangle\rangle}(b^{\dagger}_{i}b_{j}+H.c.)+V_{1}\sum_{\langle i,j\rangle}n_{i}n_{j}-\mu\sum_{i}n_{i},$ (8)
where $b^{\dagger}_{i}$ and $b_{i}$ are bosonic creation and annihilation operators and $n_{i}=b^{\dagger}_{i}b_{i}$ is the number operator at a site $i$. In this description, $\langle i,j\rangle$ and $\langle\langle i,j\rangle\rangle$ represent nearest-neighbor (NN) and next-nearest-neighbor (NNN) site pairs respectively. The hopping amplitude between NNN sites is denoted by $t_{2}$, the NN repulsion between the HCBs is represented by $V_{1}$ and $\mu$ indicates the chemical potential.
In Ref. [25] this model was thoroughly investigated using SSE QMC technique by calculating two order parameters, structure factor and superfluid density. The structure factor per site is given as,
$S(\boldsymbol{Q})=\frac{1}{N_{s}^{2}}\sum_{i,j}e^{i\boldsymbol{Q}\cdot(r_{i}-r_{j})}\langle n_{i}n_{j}\rangle,$ (9)
[…]
$\rho_{s}=\frac{1}{2\beta}\langle W_{x}^{2}+W_{y}^{2}\rangle.$ (10)
Figure 3: Top: ML prediction for the two phases, cS and SF, as a function of the NN repulsion $V_{1}$ for a fixed particle density $\rho=1/2$. Bottom: QMC results for the observables, superfluid density $\rho_{s}$ and structure factor $S(\pi,\pi)$ as a function of $V_{1}$ for a $12\times 12$ square lattice at half-filling.
Figure 4: Top: ML prediction for the three phases, cS, cSS and MI, as a function of the chemical potential $\mu$. This is for a fixed value of NN repulsion $V_{1}=4$. The vertical dashed lines indicate the ML predicted phase boundaries. Bottom: QMC results for the observables, particle-density $\rho$, superfluid density $\rho_{s}$ and structure factor $S(\pi,\pi)$ with varying chemical potential $\mu$, on a $12\times 12$ square lattice. The left $y$-axis shows the variation of $\rho_{s}$ and $S(\pi,\pi)$, while the right $y$-axis represents the variation of $\rho$.
The study revealed that at half-filling, as the NN repulsion is increased beyond a critical value $V_{c}$, the system undergoes a mixed-order phase transition (displaying the extreme Thouless effect) from a superfluid (SF) phase to a checkerboard solid (cS) phase, where either one of the sub-lattices is completely occupied. For $V>V_{c}$ as the chemical potential $\mu$ is varied, the system goes through two continuous phase transitions. The first transition takes the system from the cS phase to a checkerboard supersolid (cSS) phase, where all the particles occupy a single sub-lattice and hop within it. At the second phase transition, the cSS phase becomes a Mott insulator (MI) at particle-density $\rho=1$, where all the lattice sites are completely filled. At intermediate particle densities between $\rho=1/2$ and $1$, the model displays either a mixed-order or continuous phase transition (separated by a tricritical point) between the SF and cSS phases as a function of $V_{1}$. In Fig. 2 the above discussion is summarized in the phase diagram of the model, obtained from the SSE QMC study on a $24\times 24$ square lattice, depicted in terms of particle-density $\rho$ and NN repulsion $V_{1}$. It should be noted that in the SSE QMC simulations, different values of particle-densities $\rho$ are accessed by tuning the chemical potential $\mu$ of the system. In particular, for the insulating phases, cS and MI, which are present at $\rho=1/2$ and $1$ respectively, one obtains plateaus in the particle-density $\rho$ as a function of $\mu$.
We apply a supervised ML model trained on the QMC spin-opstring data to predict the phase transitions. For the training, we collect $10000$ spin-opstring data in a small window from deep within each of the phases : cS, cSS, SF and MI [^1] In particular for cS and MI phases, deep within the phase precisely means the $\mu$ values around center of the plateau in $\rho$.. The trained ML model is then used to predict unlabeled data along two lines in the phase diagram: firstly, the vertical line at half-filling, where the system undergoes a mixed-order phase transition from SF to cS as a function of $V_{1}$, and secondly, the horizontal line where at a fixed $V_{1}=4.0$ as function of $\mu$, and hence as function of particle-density $\rho$, the system goes through two continuous transitions, from cS to cSS and then cSS to a MI at $\rho=1$. These two lines are indicated by shaded grey bands in Fig. 2. It should be noted that we observe considerable variation in the opstring lengths among different phases. The QMC simulation in MI phase generates the largest opstring lengths, whereas in the SF phase it is the smallest.
Figure 5: Comparison of superfluid density obtained from the ML regression model with the QMC results obtained for a $12\times 12$ square lattice. The error band of the ML data has been obtained from 100 independently trained ML models.
In the top figure of Fig. 3, we show the ML prediction when tested on a dataset obtained for a $12\times 12$ square lattice at half-filling with a varying NN repulsion $V_{1}$. The figure clearly demonstrates a sharp transition between the SF and cS phase displayed by a jump between the extremum probability values of the output neurons. We contrast the ML prediction with the QMC results on the same lattice size, shown in the bottom of Fig. 3, where two observables, superfluid density $\rho_{s}$ and structure factor $S(\pi,\pi)$, are plotted with increasing NN repulsion. For $V_{1}$ below a critical value, the plot indicates the presence of the SF phase, characterized by a near-maximal value of superfluid density and a zero structure factor. In contrast, for $V_{1}$ exceeding this critical value, the system transitions to the cS phase, where $S(\pi,\pi)$ reaches its maximum value, and $\rho_{s}$ drops to zero. This sharp jump in the order parameters to their respective extrema values is a signature of the extreme Thouless effect and this behavior is also reflected in the ML predictions. Notably, the transition point predicted by the ML model matches perfectly with the QMC result, identifying the critical NN repulsion as $V_{1}=2.36\pm 0.04$. It is important to mention here that we found the Batch Normalization layer to be a crucial ingredient in the ML model for making consistent predictions. We provide details regarding this in Appendix A.
Next, in the top panel of Fig. 4, we present the ML predictions for the same system size at fixed NN repulsion $V_{1}=4.0$ as a function of the chemical potential $\mu$. The plot shows the existence of two phase transitions which appear to be continuous, the first one between phases cS and cSS and the second one between cSS and MI. In the bottom figure, we show the same phase transitions demonstrated by the QMC observables $\rho_{s}$ and $S(\pi,\pi)$, along with the particle-density $\rho$ as a function of $\mu$. The figure illustrates the existence of three phases: checkerboard solid at density $\rho=1/2$ where only the structure factor $S(\pi,\pi)$ is non-zero, checkerboard supersolid for intermediate densities where both $S(\pi,\pi)$ and $\rho_{s}$ are non-zero and the Mott insulator at density $\rho=1$ where both the observables, $S(\pi,\pi)$ and $\rho_{s}$, are zero. In the phase diagram in Fig. 2, these two transitions are depicted in terms of the corresponding particle density $\rho$ highlighted by the horizontal shaded band. In this case as well the ML predictions match well with the QMC results.
Finally, we wanted to see how much information can ML extract from our proposed input type in regression tasks. For this we selected the observable superfluid density $\rho_{s}$, which is connected to the winding of particles through the spatial dimensions of the system, thereby capturing non-local properties of the quantum system. The regression task is performed along the same line in the phase diagram as in Fig. 4, with a fixed NN repulsion $V_{1}=4.0$. To train the model we used the same data collected as before, but this time the values of $\rho_{s}$ are used as labels. Since $\rho_{s}$ is an expectation value calculated by averaging over several simulation cells, we have labeled all the spin-opstring data at a given $\mu$ with the same value, i.e., the expectation value $\rho_{s}$ at that $\mu$. Similarly, ML predictions are made for multiple spin-opstring data at a fixed value of $\mu$, and the results are then averaged to obtain the final predicted $\rho_{s}$. In Fig. 5, we compare the prediction of the regression model with the QMC results obtained for a $12\times 12$ square lattice. For the most part the predicted superfluid density values closely agree with the QMC results. However, near the transition points unsurprisingly the machine prediction show deviations as it encounters critical behavior which is absent in the training data. The regression results further support the suitability of our proposed input data for machine learning applications.
Figure 6: Phase diagram of 1D SSH model in terms of hopping amplitudes $v$ and $w$. The shaded gray band represents the line along which we apply our ML model to predict the phase transition.
Figure 7: ML prediction of the SSH model performed on lattice with $32$ sites. As $v$ is varied from 00 to $1$, $w$ changes according to $1-v$. The dashed line represents the phase boundary predicted by the ML model.
IV.2 One-dimensional Su-Schrieffer-Heeger (SSH) model
In the previous section we evaluated our input data on the $t_{2}-V_{1}$ model which features conventional phase transitions driven by symmetry breaking and defined by order parameters. However, it would be intriguing to explore how this spin-opstring input data performs in the context of topological phase transitions. Unlike conventional transitions, topological transitions do not involve any local order parameter, but are instead characterized by changes in global topological properties of the system, such as the winding number or Chern number.
Figure 8: Variance $\lambda_{1}$ of the first principal component obtained from PCA done on data collected from deep within the empty phase ($\rho=0$) and Mott insulator (MI) phase. The top plot shows the result of the analysis on the combined spin-opstring data while the bottom plot shows the case with only opstring data.
In order to demonstrate the performance of the spin-opstring input data on topological phase transitions, we select the simplest and most fundamental model for topological excitations: the one-dimensional Su-Schrieffer-Heeger (SSH) model. Originally, the SSH model was proposed to describe the behavior of electrons and the emergence of solitonic defects in a polyacetylene molecule [26]. It is a tight-binding model that describes spinless electrons on a one-dimensional dimerized chain, arising from alternating strong and weak hopping amplitudes between the nearest-neighbor sites. The Hamiltonian of the 1D SSH model is described as,
[…]
where $c^{\dagger}$ and $c$ denote the electron creation and annihilation operators, respectively. The summations in the first and second term span over the odd and even lattice sites, respectively. The parameters $v$ and $w$ correspond to the hopping amplitudes along alternate bonds. As a result of the competition between these hopping amplitudes, the SSH model undergoes a topological phase transition at half-filling. A non-topological dimer phase ($v>w$) transitions into a topological phase ($v<w$) at the critical point $v=w$, for a fixed choice of boundary. The phase diagram of 1D SSH model in terms of the hopping amplitudes $v$ and $w$, are presented in Fig. 6.
To train the ML model for 1D SSH model, we collect the spin-opstring data in a small window around two points $(v,w)=(0.9,0.1)$ and $(0.1,0.9)$ which are symmetrically apart from the phase transition at $(0.5,0.5)$. These data are labeled “topological” and “non-topological” respectively in the training process. The trained model is now fed unseen spin-opstring data along a line in the coupling space characterized by the equation $v=1-w$ for $v\in[0,1]$ which goes through the above-mentioned phase transition point. This line is indicated by the oblique shaded band in the phase diagram in Fig. 6. The predictions of the ML model is shown in Fig. 7, where we can see that it successfully predicts the true phase transition at $(0.5,0.5)$ based on the spin-opstring data. This demonstrates the spin-opstring input data’s capability to capture topological phase transitions that are independent of any local order parameter.
Figure 9: Variance $\lambda_{1}$ along the first principal component obtained from PCA done on spin-opstring (Left) and spinstate (Right) data independently for values of coupling across the transition in the $1D$ SSH model.
IV.3 Spin-state or opstring ?
We have seen above that the combination of initial spin-state and opstring obtained from QMC simulations can be successfully used for classification and regression ML tasks. It would be interesting to ask whether only the spinstate or opstring may be used without the other for training and prediction. The operator string combined with a spin configuration of the lattice is equivalent to an entire QMC simulation cell which embodies quantum features unique to a particular phase. However, we find that even when trained solely on the opstring data a ML model is capable of predicting the phase transitions of the $2D$ $t_{2}-V_{1}$ model, discussed above in Sec. IV.1. But, the opstring may not be effective as a standalone training input between two phases which are related by some symmetry. For example, in the $2D$ $t_{2}-V_{1}$ model, the empty phase at $\rho=0$ and the Mott insulator at $\rho=1$, which are related by a particle-hole symmetry transformation, have similar QMC operator strings but very different spin-states. To illustrate this, in Fig. 8 we show the variance along first principal component from a principal component analysis (PCA) performed separately on spin-opstring and opstring data, collected deep within the empty and the MI phases. While the distinction between the combined spin-opstring in the two phases is evident, the operator strings exhibit similar features in both. This demonstrates that the opstring data alone may not always be sufficient for successful ML training.
It is well-known by now that spin configurations can be used to predict thermal phase transitions with great accuracy in classical spin systems. However, for quantum phase transitions, additional information beyond the spin state is typically needed. This is clearly illustrated in Fig. 9 for the $1D$ SSH model, where we show the results of a PCA performed separately on the spin-state and the combined spin-opstring data across the transition between the topological and non-topological phases. The principal components (the variance of first one shown in Fig. 9) for the spin-state data fail to reveal any distinct features across a wide range of couplings during the transition, whereas the spin-opstring data exhibits a clear trend from which even the phase transition can be determined (by the location of the trough). In conclusion, while either the spin-state or opstring data may serve as an effective ML input in some cases, combining both ensures more robust and reliable results in all scenarios.
V Conclusions
In conclusion, based on Stochastic Series Expansion QMC we have proposed a new input data type “spin-opstring” for machine learning applications. The spin-opstring is a compact representation of the QMC simulation cell that retains all quantum information of the system while being memory-efficient. It consists of the initial state followed by an operator string that describes the state’s evolution through imaginary time, ultimately bringing it back to its original configuration. In order to demonstrate the applicability of spin-opstring as ML input, we have applied supervised machine learning on two different models : the two-dimensional $t_{2}-V_{1}$ model and the one-dimensional Su-Schrieffer-Heeger model. We found that in both models the input data performs very well by capturing the phase transitions accurately. This highlights the versatility of our input data in identifying conventional phase transitions, both continuous and discontinuous, as well as topological phase transitions. Moreover, we employed the spin-opstring to perform a regression task predicting the superfluid density in the $2D$ $t_{2}-V_{1}$ model. These results provides additional evidence for the suitability of our proposed input data in machine learning applications. It would be interesting to see how the spin-opstring data perform in the domain of unsupervised ML algorithms.
Acknowledgements.
This work was supported by the Center for Theory and Computation, National Tsing Hua University-113H0001L9, Hsinchu 300 Taiwan, the Taiwan National Science and Technology Council Project No. 112-2639-M-002-006-ASP and NSTC 113-2112-M-002-033-MY3.
Appendix A Effect of Batch Normalization
Batch Normalization (BN) is a well-known method in neural network machine learning which helps to train the models faster and also acts as a regularization to prevent overfitting [30]. It acts by first“whitening” i.e., transforming linearly to have zero mean and unit variance, the outputs of a network layer and further scale and shift the resulting output with trainable parameters. This results in better convergence of the gradient descent training process. In our study of phase classification, we have found that BN plays an important role in stabilizing the predictions of the phase transition points.
In Fig. 10, we compare the machine predictions for the SF-cS transition in the 2D $t_{2}-V_{1}$ model at half-filling, as a function of NN repulsion $V_{1}$, obtained by training the same network with and without BN layers. We show the predictions for $100$ independently trained models in each case using the same hyperparameters. The figure clearly shows that the ML model gives consistent predictions for the phase transition point when BN is applied. Although the predictions are correct deep in the phases, the predictions without BN show a wide variation in the predicted phase transition point.
Figure 10: Effect of Batch Normalization (BN) on machine prediction for the SF-cS transition in the 2D $t_{2}-V_{1}$ model at half-filling, as a function of NN repulsion $V_{1}$. Top: Predictions from $100$ independently trained models without BN. Bottom: The same number of independent predictions from the ML model with added BN layers.